OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (anonymous):

HELPPPP

OpenStudy (zzr0ck3r):

y = mx + b system (-2,3)(2,7) 3 = -2m + b 7 = 2m + b add them together 10=2b b = 5 7 = 2m+5 m = 1 so y = x+5 or x-y = -5 make since how I did that? you could have done substitution also but this was easier imo

OpenStudy (anonymous):

so its C

OpenStudy (zzr0ck3r):

if c is the third one

OpenStudy (zzr0ck3r):

but do you see how I did it?

OpenStudy (anonymous):

yea thanks could you help with this too. If A = {negative even integers} and B = {negative odd integers}, what is A B? {0} {negative integers} {integers} { }

OpenStudy (zzr0ck3r):

pos even integer

OpenStudy (zzr0ck3r):

-2*-3 = 6 -4*-7 = 28...

OpenStudy (anonymous):

so wait which one is it

OpenStudy (zzr0ck3r):

I guess c but its odd they say all this stuff about even odd and only give pos integers. they could have just asked what a neg*neg is

OpenStudy (zzr0ck3r):

is the last option just an empty set?

OpenStudy (anonymous):

how bout this .. Connie needs to burn at least 300 calories at the gym today. She can burn 20 calories per minute on the treadmill and 10 calories per minute on the elliptical trainer. Which inequality below can be used to find the possible combinations of time she could spend on each machine and still meet her goal of burning at least 300 calories? 20t + 10e =>300 20t - 10e =>300 20t + 10e< = 300 20t - 10e<= 300